OpenStudy (anonymous):

Write an equation for the quadratic function whose graph contains the given vertex and point. Vertex: (5,-2) and point: (-4, 241)

OpenStudy (amoodarya):

(x-a)^2=4p(y-b) (a,b)=vertex=(5,-2) so (x-5)^2=4p(y+2) now put (-4,241) in it and find p so you have equation

OpenStudy (amoodarya):

another solution is y=a(x-5)^2+b now put (-4,241) , (5,-2) in it and find a,b